OpenStudy (anonymous):

Polynomial Operations... SIMPLIFY: (4y^2-5y) + (3y-7y^2) - (2y^2+6y-5) A.)9y^2-8y-5 C.)-3y^2+2y+5 B.)-5y^2-8y+5 D.)5y^2-6y-5

OpenStudy (anonymous):

I can eliminate A and B

OpenStudy (anonymous):

From Mathematica:\[(4*y{}^{\wedge}2 - 5*y) + (3*y - 7*y{}^{\wedge}2) - (2*y{}^{\wedge}2 + 6*y - 5)\text{ = }5-8 y-5 y{}^{\wedge}2 \]

OpenStudy (anonymous):

^looks better than mine

jimthompson5910 (jim_thompson5910):

not sure how you eliminated choice B when choice B is the correct answer.

OpenStudy (anonymous):

\[(4*y{}^{\wedge}2 - 5*y) + (3*y - 7*y{}^{\wedge}2) - (2*y{}^{\wedge}2 + 6*y - 5) \]\[\left(-5 y+4 y^2\right) + \left(3 y-7 y^2\right) - \left(-5+6 y+2 y^2\right) \]\[-2 y-3 y^2 - \left(-5+6 y+2 y^2\right) \]\[-5 y^2-8 y+5 \]B. is the answer

OpenStudy (anonymous):

@jim_thompson5910 Thanks...

OpenStudy (anonymous):

@robtobey Thks for explainng

OpenStudy (anonymous):

You are welcome.